# Excited $\Omega$ hyperon in charmful $\Omega_{b}$ weak decays

### v1

@@ I Introduction
In the above equation, $J/\psi(\textbf{q};1,J_{J/\psi}^{z})$, $\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})$, and $\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})$ are wave functions, where ${\bf q}$ and ${\bf P}_{i(f)}$ are the total momentum, $J_{i(f)}$ the total angular momentum, and $J^{z}_{i(f)}$ the third component of the total angular momentum.
To calculate $\mathcal{M}^{PC(PV)}$ in Eq. (II), the details of the wave functions are required. Within the non-relativistic constituent quark model, the spatial wave function of a baryon consisting of $q_{1}q_{2}q_{3}$ can be represented as a product of the $\rho$-oscillator part and the $\lambda$-oscillator part, where $\rho$ corresponds to the $q_{1}-q_{2}$ oscillation and $\lambda$ corresponds to the $\rho-q_{3}$ oscillation. Hence, the internal momenta of the $\rho$- and $\lambda$-oscillators $\textbf{p}_{\rho}$ and $\textbf{p}_{\lambda}$ are expressed as
@@
$\mathbf{p}_{\lambda}$ $=$ $\frac{\sqrt{6}}{2}\frac{m_{3}(\textbf{p}_{1}+\textbf{p}_{2})-(m_{1}+m_{2})\textbf{p}_{3}}{m_{1}+m_{2}+m_{3}}\,,$ (8)
@@
In momentum space, we present the baryon wave function as Liu:2019wdr
$\Psi^{\sigma}_{NLM_{L}}(\textbf{p}_{\rho},\textbf{p}_{\lambda})\!=\!\!\!\sum_{N,M_{L}}C^{n_{\rho}l_{\rho}m_{\rho}}_{n_{\lambda}l_{\lambda}m_{\lambda}}\left[\psi_{n_{\rho}l_{\rho}m_{\rho}}(\textbf{p}_{\rho})\psi_{n_{\lambda}l_{\lambda}m_{\lambda}}(\textbf{p}_{\lambda})\right]^{\sigma}_{NLM_{L}},$ (9)
where the superscript $\sigma$ characterizes the same set of quantum numbers $(N,L,M_{L})$ arising from different combinations of the $\rho$ and $\lambda$ oscillation systems Karl:1969iz ; Zhenping:1991 ; Xiao:2013xi . Specifically, the combinations are defined by $N=2(n_{\rho}+n_{\lambda})+l_{\rho}+l_{\lambda}$, $L=l_{\rho}+l_{\lambda}$, and $M_{L}=m_{\rho}+m_{\lambda}$, where $(n_{i},l_{i},m_{i}$) with $i=\rho$ or $\lambda$ are the principal, orbital, and magnetic quantum numbers, respectively. In the $N=1$ and 2 shells, the $1S$-, $1P$-, and $1D$-wave, as well as the $2S$-wave $\Omega$ wave functions, have been studied in Ref. Liu:2019wdr within the harmonic oscillator model, providing the information on the coefficient $C^{n_{\rho}l_{\rho}m_{\rho}}_{n_{\lambda}l_{\lambda}m_{\lambda}}$.
In Eq. (9), $\psi_{nlm}(\textbf{p})$ is expressed by
@@
where the $l$th solid harmonic polynomial is defined as $\mathcal{Y}_{lm}(\textbf{p})=|\textbf{p}|^{l}Y_{lm}(\mathbf{\hat{p}})$, and the oscillator parameter $\alpha$ can be either $\alpha_{\rho}$ or $\alpha_{\lambda}$. For the $\Omega_{b}^{-}$ wave function, the parameter $\alpha_{\lambda}$ is given by $\alpha_{\lambda}=[(3m_{b})(2m_{s}+m_{b})]^{1/4}\alpha_{\rho}$. For the $\Omega^{(*)-}$ wave function, we define $\alpha\equiv\alpha_{\rho}=\alpha_{\lambda}$. The flavor and spin wave functions of the $\Omega_{b}^{-}$ and $\Omega^{(*)-}$ baryons should be incorporated into the spatial wave functions. These can be found in Refs. Xiao:2013xi ; Wang:2017kfr ; Yao:2018jmc ; Zhong:2007gp .
In the $J/\psi$ wave function:
$\Psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})=\varphi_{J/\psi}\chi^{1}_{s_{z}}\psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})\,,$ (11)
$\varphi_{J/\psi}$, $\chi^{1}_{s_{z}}$, and $\psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})$ represent the flavor, spin, and spatial wave functions, respectively, which are given as:
$\chi^{1}_{1,0,-1}=(\uparrow\uparrow,(\uparrow\downarrow+\downarrow\uparrow)/\sqrt{2},\downarrow\downarrow)\,,$
$\varphi_{J/\psi}=(c\bar{c}-\bar{c}c)/\sqrt{2}\,,$
$\psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})=\frac{1}{\pi^{3/4}\beta^{3/2}}\mathrm{exp}\left[-\frac{(\textbf{p}^{\prime}_{3}-\textbf{p}_{4})^{2}}{8\beta^{2}}\right]\,.$ (12)
In Eq. (II), $s_{z}$ denotes the $z$ component of the $J/\psi$ spin. The term $\varphi_{J/\psi}$ respects $C$-parity, while $\beta$ in $\psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})$ controls the width of Gaussian distribution in momentum space.
To convert ${\cal M}(\Omega_{b}^{-}\to J/\psi\Omega^{(*)-})$ from Eqs. (6) and (II) into the decay width, we use the following equation:
$\Gamma=8\pi^{2}\frac{|\textbf{q}|E_{J/\psi}E_{\Omega^{(*)}}}{M_{\Omega_{b}}}\frac{1}{2J_{\Omega_{b}}+1}\sum_{\texttt{spins}}|\mathcal{M}(\Omega_{b}^{-}\to J/\psi\Omega^{(*)-})|^{2}\,,$ (13)
which is applied in our numerical analysis.
@@ III Numerical results
For our numerical analysis, we adopt the CKM matrix elements and the masses of $\Omega_{(b)}^{-}$ and $J/\psi$ from PDG pdg , while the quark masses are taken from Wang:2017kfr , as follows:
@@
The lifetime of the $\Omega_{b}^{-}$ state is also taken from PDG pdg : $\tau_{\Omega_{b}}=1.64\times 10^{-12}s$. The Wilson coefficient $C_{2}=-0.365$ is from Hsiao:2021mlp ; Ali:1998eb ; Hsiao:2014mua . For the harmonic oscillator parameters, we use $(\alpha_{\lambda},\alpha_{\rho})=(0.56,0.44)$ GeV for $\Omega_{b}^{-}$ Zhong:2007gp , and $\beta=0.50$ GeV for $J/\psi$, as adopted in Barnes:2005pb ; Xiao:2018iez . For the ground-state and excited $\Omega$ hyperon states, the masses and the oscillator parameters $\alpha$, calculated in Liu:2019wdr , are summarized in Table 1.
Table 1: Our results for $\Omega_{b}^{-}\to J/\psi\Omega^{(*)-}$ using the constituent model, with the quantum numbers $n^{2S+1}J^{P}$ assigned to $\Omega^{(*)-}$. The parameter $\alpha$ and the mass of $\Omega^{(*)-}$ $M_{f}$ are both given in units of MeV; $\Gamma$ and ${\cal B}$ are in units of $10^{-17}$ GeV and $10^{-4}$, respectively.
@@
$\Omega(1^{4}S_{\frac{3}{2}^{+}})$ 440 1672 8.3 2.1 1.0
$\Omega(1^{2}P_{\frac{1}{2}^{-}})$ 428 1957 4.5 1.1 0.52
$\Omega(1^{2}P_{\frac{3}{2}^{-}})$ 411 2012 10.7 2.7 1.2
$\Omega(2^{2}S_{\frac{1}{2}^{+}})$ 387 2232 5.1 1.2 0.57
$\Omega(2^{4}S_{\frac{3}{2}^{+}})$ 381 2159 20.6 5.1 2.4
$\Omega(1^{2}D_{\frac{3}{2}^{+}})$ 394 2245 2.3 0.58 0.27
$\Omega(1^{2}D_{\frac{5}{2}^{+}})$ 380 2303 5.1 1.2 0.57
$\Omega(1^{4}D_{\frac{1}{2}^{+}})$ 413 2141 2.6 0.64 0.30
$\Omega(1^{4}D_{\frac{3}{2}^{+}})$ 399 2188 6.0 1.4 0.67
$\Omega(1^{4}D_{\frac{5}{2}^{+}})$ 383 2252 8.2 2.1 1.0
$\Omega(1^{4}D_{\frac{7}{2}^{+}})$ 367 2321 4.2 1.0 0.47
Utilizing the inputs above, we calculate ${\cal B}(\Omega_{b}^{-}\to J/\psi\Omega^{-})$ to compare it with the previous studies. Specifically, the decay widths and branching fractions of possible $\Omega_{b}^{-}\to J/\psi\Omega^{(*)-}$ are presented in Table 1. Additionally, we plot $\Gamma(\Omega_{b}^{-}\to J/\psi\Omega^{-})$ as functions of the parameters $\alpha$ and $\beta$ in Fig. 2 and $\Gamma(\Omega_{b}^{-}\to J/\psi\Omega^{*-})$ as a function of $m_{\Omega^{*}}$ in Fig. 3.
Table 2: The branching fraction of $\Omega_{b}^{0}\rightarrow J/\psi\Omega^{-}$ is expressed in units of $10^{-4}$, compared with results from different models.
Our work Ref. Gutsche:2018utw Ref. Fayyazuddin:2017sxq Ref. Hsiao:2021mlp Ref. Cheng:1996cs Ref. Rui:2023fiz
@@
Figure 2: The decay width of $\Omega_{b}^{-}\to J/\psi\Omega^{-}$ varies with the harmonic parameters $\alpha$ of $\Omega$ baryon and $\beta$ of $J/\psi$ meson.

### v4

@@ I Introduction
Here, the factor $\frac{1}{\sqrt{3}}$ is the normalization coefficient and $\hat{P}_{ij}$ represents the permutation operator of quarks $i$ and $j$. In the above equation, $J/\psi(\textbf{q};1,J_{J/\psi}^{z})$, $\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})$, and $\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})$ are wave functions, where ${\bf q}$ and ${\bf P}_{i(f)}$ are the total momentum, $J_{i(f)}$ are the total angular momentum, and $J^{z}_{i(f)}$ are the third component of the total angular momentum.
It is worth noting that the calculation of the amplitudes for $\Omega_{b}\to\Omega^{(*)}J/\psi$ decays should consider a question of symmetry. The flavor wave functions of $\Omega_{b}$ and $\Omega$ are $\varphi_{\Omega_{b}}=ssb$ and $\varphi_{\Omega}=sss$, respectively. In the $\Omega_{b}\rightarrow\Omega^{(*)}J/\psi$ nonleptonic decay process, the bottom quark of the $\Omega_{b}$ decays into an $s$ quark, which can be any of the three $s$ quarks in the $\Omega$. Since the wave functions of the final state $\Omega^{(*)}$ baryon are fully antisymmetric under the exchange of any two quarks, so we obtain
$\langle J/\psi(\textbf{q};1,J_{J/\psi}^{z})\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})|H_{W}^{PC(PV)}|\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})\rangle$
$=$ $\langle J/\psi(\textbf{q};1,J_{J/\psi}^{z})\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})|\hat{P}_{25}H_{W}^{PC(PV)}|\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})\rangle$
$=$ $\langle J/\psi(\textbf{q};1,J_{J/\psi}^{z})\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})|\hat{P}_{15}H_{W}^{PC(PV)}|\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})\rangle.$
Therefore, one can obtain
$\mathcal{M}^{PC(PV)}$ $=$ $\sqrt{3}\langle J/\psi(\textbf{q};1,J_{J/\psi}^{z})\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})|$ (20)
$H_{W}^{PC(PV)}|\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})\rangle.$
Similar calculations have been applied in Ref. Pervin:2006ie .
Table 1: The spin-flavor-space wave-functions of $\Omega$ baryons under $SU(6)$ quark model classification are listed below. We denote the baryon states as $|N_{6},^{2S+1}N_{3},N,L,J^{P}\rangle$ where $N_{6}$ stands for the irreducible representation of spin-flavor $SU(6)$ group, $N_{3}$ stands for the irreducible representation of flavor $SU(3)$ group and $N$, $L$, $J^{P}$ as principal quantum number, total orbital angular momentum and spin-parity, respectivelyXiao:2018pwe ; Liu:2019wdr ; Wang:2018hmi . The $\phi$, $\chi$, $\psi$ denote flavor, spin and spatial wave function, respectively. The Clebsch-Gorden coefficients of spin-orbital coupling and color wave function $\zeta_{\Omega}$ have been omitted.
$n^{2S+1}L_{J^{p}}$ $|N_{6},^{2S+1}N_{3},N,L,J^{P}\rangle$ Wave function
$1^{4}S_{\frac{3}{2}^{+}}$ $|56,^{4}10,0,0,\frac{3}{2}^{+}\rangle$ $\phi^{s}\chi^{s}\psi_{000}^{s}$
$1^{2}P_{\frac{1}{2}^{-}}$ $|70,^{2}10,1,1,\frac{1}{2}^{-}\rangle$ $\frac{1}{\sqrt{2}}\left(\phi^{s}\chi^{\rho}\psi_{11L_{z}}^{\rho}+\phi^{s}\chi^{\lambda}\psi_{11L_{z}}^{\lambda}\right)$
$1^{2}P_{\frac{3}{2}^{-}}$ $|70,^{2}10,1,1,\frac{3}{2}^{-}\rangle$ $\frac{1}{\sqrt{2}}\left(\phi^{s}\chi^{\rho}\psi_{11L_{z}}^{\rho}+\phi^{s}\chi^{\lambda}\psi_{11L_{z}}^{\lambda}\right)$
$2^{2}S_{\frac{1}{2}^{+}}$ $|70,^{2}10,2,0,\frac{1}{2}^{+}\rangle$ $\frac{1}{\sqrt{2}}\left(\phi^{s}\chi^{\rho}\psi_{200}^{\rho}+\phi^{s}\chi^{\lambda}\psi_{200}^{\lambda}\right)$
$2^{4}S_{\frac{3}{2}^{+}}$ $|56,^{4}10,2,0,\frac{3}{2}^{+}\rangle$ $\phi^{s}\chi^{s}\psi_{200}^{s}$
$1^{2}D_{\frac{3}{2}^{+}}$ $|70,^{2}10,2,2,\frac{3}{2}^{+}\rangle$ $\frac{1}{\sqrt{2}}\left(\phi^{s}\chi^{\rho}\psi_{22L_{z}}^{\rho}+\phi^{s}\chi^{\lambda}\psi_{22L_{z}}^{\lambda}\right)$
$1^{2}D_{\frac{5}{2}^{+}}$ $|70,^{2}10,2,2,\frac{5}{2}^{+}\rangle$ $\frac{1}{\sqrt{2}}\left(\phi^{s}\chi^{\rho}\psi_{22L_{z}}^{\rho}+\phi^{s}\chi^{\lambda}\psi_{22L_{z}}^{\lambda}\right)$
$1^{4}D_{\frac{1}{2}^{+}}$ $|56,^{4}10,2,2,\frac{1}{2}^{+}\rangle$ $\phi^{s}\chi^{s}\psi_{22L_{z}}^{s}$
$1^{4}D_{\frac{3}{2}^{+}}$ $|56,^{4}10,2,2,\frac{3}{2}^{+}\rangle$ $\phi^{s}\chi^{s}\psi_{22L_{z}}^{s}$
$1^{4}D_{\frac{5}{2}^{+}}$ $|56,^{4}10,2,2,\frac{5}{2}^{+}\rangle$ $\phi^{s}\chi^{s}\psi_{22L_{z}}^{s}$
$1^{4}D_{\frac{7}{2}^{+}}$ $|56,^{4}10,2,2,\frac{7}{2}^{+}\rangle$ $\phi^{s}\chi^{s}\psi_{22L_{z}}^{s}$
Table 2: The spatial functions $\psi^{\sigma}_{NLM_{L}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})$ as the linear combination of $\psi_{n_{\rho}l_{\rho}m_{\rho}}(\mathbf{p}_{\rho})\psi_{n_{\lambda}l_{\lambda}m_{\lambda}}(\mathbf{p}_{\lambda})$.
$\psi^{S}_{000}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\psi_{000}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})$
$\psi^{\rho}_{11M_{L}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\psi_{01M_{L}}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})$
$\psi^{\lambda}_{11M_{L}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\psi_{000}(\mathbf{p}_{\rho})\psi_{01M_{L}}(\mathbf{p}_{\lambda})$
$\psi^{S}_{200}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\frac{1}{\sqrt{2}}[\psi_{100}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})+\psi_{000}(\mathbf{p}_{\rho})\psi_{100}(\mathbf{p}_{\lambda})]$
$\psi^{\lambda}_{200}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\frac{1}{\sqrt{2}}[-\psi_{100}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})+\psi_{000}(\mathbf{p}_{\rho})\psi_{100}(\mathbf{p}_{\lambda})]$
$\psi^{\rho}_{200}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})$
$=\frac{1}{\sqrt{3}}[\psi_{011}(\mathbf{p}_{\rho})\psi_{01-1}(\mathbf{p}_{\lambda})-\psi_{010}(\mathbf{p}_{\rho})\psi_{010}(\mathbf{p}_{\lambda})+\psi_{01-1}(\mathbf{p}_{\rho})\psi_{011}(\mathbf{p}_{\lambda})]$
$\psi^{S}_{22M_{L}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\frac{1}{\sqrt{2}}[\psi_{02M_{L}}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})+\psi_{000}(\mathbf{p}_{\rho})\psi_{02M_{L}}(\mathbf{p}_{\lambda})]$
$\psi^{\lambda}_{22M_{L}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\frac{1}{\sqrt{2}}[\psi_{02M_{L}}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})-\psi_{000}(\mathbf{p}_{\rho})\psi_{02M_{L}}(\mathbf{p}_{\lambda})]$
$\psi^{\rho}_{222}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\psi_{011}(\mathbf{p}_{\rho})\psi_{011}(\mathbf{p}_{\lambda})$
$\psi^{\rho}_{221}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\frac{1}{\sqrt{2}}[\psi_{010}(\mathbf{p}_{\rho})\psi_{011}(\mathbf{p}_{\lambda})+\psi_{011}(\mathbf{p}_{\rho})\psi_{010}(\mathbf{p}_{\lambda})]$
$\psi^{\rho}_{220}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})$
$=\frac{1}{\sqrt{6}}[\psi_{01-1}(\mathbf{p}_{\rho})\psi_{011}(\mathbf{p}_{\lambda})+2\psi_{010}(\mathbf{p}_{\rho})\psi_{010}(\mathbf{p}_{\lambda})+\psi_{011}(\mathbf{p}_{\rho})\psi_{01-1}(\mathbf{p}_{\lambda})]$
$\psi^{\rho}_{22-1}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\frac{1}{\sqrt{2}}[\psi_{01-1}(\mathbf{p}_{\rho})\psi_{010}(\mathbf{p}_{\lambda})+\psi_{010}(\mathbf{p}_{\rho})\psi_{01-1}(\mathbf{p}_{\lambda})]$
$\psi^{\rho}_{22-2}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\psi_{01-1}(\mathbf{p}_{\rho})\psi_{01-1}(\mathbf{p}_{\lambda})$
To calculate $\mathcal{M}^{PC(PV)}$ in Eq. (20), the details of the wave functions are required. The total wave function of a baryon system should include four parts: a color wave function $\zeta$, a flavor wave function $\phi$, a spin wave function $\chi$, and a spatial wave function $\psi$. The spin wave function of a baryon satisfies SU(2) symmetry, which can be expressed as symmetric ($\chi^{S}_{S_{z}}$), mixed antisymmetric($\chi^{\rho}_{S_{z}}$), mixed symmetric($\chi^{\lambda}_{S_{z}}$) spin wave functions. The detailed representations of the spin wave function is shown in Refs. Niu:2020gjw ; Wang:2017kfr ; Xiao:2013xi .
For the $\Omega$ spectrum, the flavor-spin wave functions are representations of SU(6), which are denoted by $|N_{6},^{2S+1}N_{3}\rangle$, where $N_{6}$ ($N_{3}$) represents the SU(6) (SU(3)) representation and $S$ stands for the total spin of the wave function, and the detailed wave function of $\Omega^{(*)}$ excited states are shown in Table 1. In momentum space, we present the simple harmonic oscillator wave functions to describe the $\Omega^{(*)}$ baryon Wang:2018hmi . The explicit forms of the spatial wave function $\Psi_{NLM_{L}}^{\sigma}(\textbf{p}_{\rho},\textbf{p}_{\lambda})$ in the momentum space, up to the $N$ = 2 shell, can be found in Table 2. The superscript $\sigma$ characterizes the same set of quantum numbers $(N,L,M_{L})$ arising from different combinations of the $\rho$ and $\lambda$ oscillation systems Karl:1969iz ; Zhenping:1991 ; Xiao:2013xi . Specifically, the combinations are defined by $N=2(n_{\rho}+n_{\lambda})+l_{\rho}+l_{\lambda}$, $L=l_{\rho}+l_{\lambda},...,|l_{\rho}-l_{\lambda}|$, and $M_{L}=m_{\rho}+m_{\lambda}$, where $(n_{i},l_{i},m_{i}$) with $i=\rho$ or $\lambda$ are the principal, orbital, and magnetic quantum numbers, respectively. The internal momenta of the $\rho$- and $\lambda$-oscillators, $\textbf{p}_{\rho}$ and $\textbf{p}_{\lambda}$, are expressed as
@@
$\mathbf{p}_{\lambda}$ $=$ $\frac{\sqrt{6}}{2}\frac{m_{5}(\textbf{p}_{1}+\textbf{p}_{2})-(m_{1}+m_{2})\textbf{p}_{5}}{m_{1}+m_{2}+m_{5}}\,,$ (21)
@@
In momentum space, $\psi_{nlm}(\textbf{p})$ is expressed by
@@
where the $l$th solid harmonic polynomial is defined as $\mathcal{Y}_{lm}(\textbf{p})=|\textbf{p}|^{l}Y_{lm}(\mathbf{\hat{p}})$, and the oscillator parameter $\alpha$ can be either $\alpha_{\rho}$ or $\alpha_{\lambda}$. For the $\Omega^{(*)}$ wave function, we define $\alpha\equiv\alpha_{\rho}=\alpha_{\lambda}$. In this work, we use the $\alpha$ value obtained from single Gaussian fitting in Ref. Liu:2019wdr .
For $\Omega_{b}$ baryon, the total wave function can be represented as
$\Psi_{\Omega_{b}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\zeta_{\Omega_{b}}\varphi_{\Omega_{b}}\chi^{\lambda}_{s_{z}}\psi_{000}^{S}(\textbf{p}_{\rho},\textbf{p}_{\lambda})\,,$ (23)
where $\zeta_{\Omega_{b}}$, $\chi^{\lambda}_{S_{z}}$, and $\psi_{000}^{S}(\textbf{p}_{\rho},\textbf{p}_{\lambda})$ represent the color, spin, and spatial wave functions, respectively. It is worth noting that in momentum space, we employ the simple harmonic oscillator wave functions to describe $\psi_{000}^{S}(\textbf{p}_{\rho},\textbf{p}_{\lambda})$, which is represented as
$\psi^{S}_{000}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\psi_{000}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})$ (24)
where $\textbf{p}_{\rho}$ and $\textbf{p}_{\lambda}$ are the internal momenta of the $\rho$- and $\lambda$-oscillators, respectively. It is worth mentioning that the parameter $\alpha_{\lambda}$ of $\Omega_{b}$ wave function is given by $\alpha_{\lambda}=[(3m_{b})/(2m_{s}+m_{b})]^{1/4}\alpha_{\rho}$ Wang:2017kfr ; Yao:2018jmc .
For the wave function of $J/\psi$ meson, which can be represented as
$\Psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})=\zeta_{J/\psi}\varphi_{J/\psi}\chi^{1}_{s_{z}}\psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})\,,$ (25)
where $\zeta_{J/\psi}$ is the color wave function. $\varphi_{J/\psi}$, $\chi^{1}_{s_{z}}$, and $\psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})$ represent the flavor, spin, and spatial wave functions, respectively, which are given as:
$\chi^{1}_{1,0,-1}=(\uparrow\uparrow,(\uparrow\downarrow+\downarrow\uparrow)/\sqrt{2},\downarrow\downarrow)\,,$ (26)
$\varphi_{J/\psi}=c\bar{c}\,,$
$\psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})=\frac{1}{\pi^{3/4}\beta^{3/2}}\mathrm{exp}\left[-\frac{(\textbf{p}^{\prime}_{3}-\textbf{p}_{4})^{2}}{8\beta^{2}}\right]\,.$
In Eq. (26), $s_{z}$ denotes the $z$ component of the $J/\psi$ spin and $\beta$ in $\psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})$ controls the width of Gaussian distribution in momentum space.
To convert ${\cal M}(\Omega_{b}\to J/\psi\Omega^{(*)})$ from Eqs. (17) and (18) into the decay width, we use the following equation:
$\Gamma=8\pi^{2}\frac{|\textbf{q}|E_{J/\psi}E_{\Omega^{(*)}}}{M_{\Omega_{b}}}\frac{1}{2J_{\Omega_{b}}+1}\sum_{J_{i}^{z},J_{f}^{z}}|\mathcal{M}_{J_{i}^{z},J_{f}^{z}}|^{2}\,,$ (27)
which is applied in our numerical analysis. In Eq. (27), $\mathcal{M}_{J_{i}^{z},J_{f}^{z}}$ is the transition amplitude.
@@ III Numerical results
For our numerical analysis, we adopt the CKM matrix elements and the masses of $\Omega_{b}$ and $J/\psi$ from PDG pdg , while the quark masses are taken from Wang:2017kfr , as follows:
@@
The lifetime of the $\Omega_{b}$ state is also taken from PDG pdg : $\tau_{\Omega_{b}}=1.64\times 10^{-12}s$. The Wilson coefficient $C_{2}=-0.365$ is from Hsiao:2021mlp ; Ali:1998eb ; Hsiao:2014mua . For the harmonic oscillator parameters, we use $(\alpha_{\lambda},\alpha_{\rho})=(0.56,0.44)$ GeV for $\Omega_{b}$ Zhong:2007gp , and $\beta=0.50$ GeV for $J/\psi$, as adopted in Barnes:2005pb ; Xiao:2018iez . For the ground-state and excited $\Omega$ hyperon states, the masses and the oscillator parameters $\alpha$, calculated in Liu:2019wdr , are summarized in Table 3.
Table 3: Our results for $\Omega_{b}\to J/\psi\Omega^{(*)}$ using the constituent model, with the quantum numbers $n^{2S+1}J^{P}$ assigned to $\Omega^{(*)}$. The parameter $\alpha$ and the mass of $\Omega^{(*)}$ $M_{f}$ are both given in units of MeV; $\Gamma$ and ${\cal B}$ are in units of $10^{-17}$ GeV and $10^{-4}$, respectively.
@@
$\Omega(1^{4}S_{\frac{3}{2}^{+}})$ 440 1672 36 8.8 1.00
$\Omega(1^{2}P_{\frac{1}{2}^{-}})$ 428 1957 17 4.2 0.48
$\Omega(1^{2}P_{\frac{3}{2}^{-}})$ 411 2012 42 11 1.25
$\Omega(2^{2}S_{\frac{1}{2}^{+}})$ 387 2232 18 4.5 0.51
$\Omega(2^{4}S_{\frac{3}{2}^{+}})$ 381 2159 79 20 2.3
$\Omega(1^{2}D_{\frac{3}{2}^{+}})$ 394 2245 7.5 1.9 0.22
$\Omega(1^{2}D_{\frac{5}{2}^{+}})$ 380 2303 17 4.2 0.48
$\Omega(1^{4}D_{\frac{1}{2}^{+}})$ 413 2141 9.3 2.3 0.26
$\Omega(1^{4}D_{\frac{3}{2}^{+}})$ 399 2188 21 5.4 0.61
$\Omega(1^{4}D_{\frac{5}{2}^{+}})$ 383 2252 29 7.2 0.81
$\Omega(1^{4}D_{\frac{7}{2}^{+}})$ 367 2321 14 3.3 0.38
Utilizing the inputs above, we calculate ${\cal B}(\Omega_{b}\to J/\psi\Omega)$ to compare it with the previous studies. Specifically, the decay widths and branching fractions of possible $\Omega_{b}\to J/\psi\Omega^{(*)}$ are presented in Table 3. Additionally, we plot $\Gamma(\Omega_{b}\to J/\psi\Omega)$ as functions of the parameters $\alpha$ and $\beta$ in Fig. 2 and $\Gamma(\Omega_{b}\to J/\psi\Omega^{*})$ as a function of $m_{\Omega^{*}}$ in Fig. 3.
Table 4: The branching fraction of $\Omega_{b}\rightarrow J/\psi\Omega$ is expressed in units of $10^{-4}$, compared with results from different models.
Our work   Gutsche:2018utw   Fayyazuddin:2017sxq   Hsiao:2021mlp   Cheng:1996cs   Rui:2023fiz   Wang:2024mjw
@@
Figure 2: The decay width of $\Omega_{b}\to J/\psi\Omega$ varies with the harmonic parameters $\alpha$ of $\Omega$ baryon and $\beta$ of $J/\psi$ meson.